# On the robustness of modeling grounded word learning through a child’s egocentric input

### v1

@@ 3.1 Vision Encoder
For embedding static video frames, we replaced the ResNeXt-CNN in CVCL with a vision transformer (ViT-B/14) (Dosovitskiy et al.,, 2020), which is separately pre-trained only using visual data from the SAYCam dataset. This pre-training step is performed separately for each child’s visual data, using the self-supervised approach from DINO (Orhan and Lake,, 2024; Caron et al.,, 2021). From the set of frames associated with a given utterance (up to 16 sampled frames), we always sample one frame randomly to be paired with the corresponding utterance. During training, data augmentation is performed by performing a random crop to each video frame, followed by resizing the frame to $224\times 224$ pixels, and occassionally horizontally flipping the frame. During evaluation, frames are instead always resized to 256 pixels along the minor edge, followed by a $224\times 224$ centered crop.
We obtain the output of the vision transformer from the [CLS] token, a single embedding of size 768. This embedding is fed through a learnable linear projection head, followed by a layer norm and a dropout layer, such that the vision encoder returns a single embedding of size $D=512$. Similar to Vong et al., (2024), the bulk of the vision encoder is frozen during training, except for the learnable linear projection head.
@@ 3.2 Variant 1: Language encoder with minimal syntax (CVCL)
In this model variant, we examine whether providing additional word order information may be useful for grounded word learning compared to a bag-of-words approach, by incorporating positional information into the embedding. Because this is a relatively minor change from the original CVCL model, we still refer to this variant as CVCL throughout.
Language Encoder (Embedding): We used the same Embedding encoder from Vong et al., (2024), which extracts separate embeddings for each word in an utterance, and computes the average across all word embeddings to produce a single utterance embedding of size $D=512$. In this variant, we add a learned, absolute positional embedding to each token embedding before the averaging step. Prior to passing utterances into the language encoder, tokenization was performed using a word-level tokenizer built via spaCy.
@@
Learning Objective (Contrastive): Following a standard approach to training vision-language models (Radford et al.,, 2021; Jia et al.,, 2021), we employ an InfoNCE contrastive learning objective, which aims to bring together frame embeddings with their corresponding utterance embeddings in a mini-batch (that is, encouraging a larger dot product between the vectors), while separating all other mismatching frame and utterance embeddings (smaller dot product between the vectors). First, we obtain visual embeddings for each frame, $\mathbf{v_{i}}=f_{\theta}(\mathbf{x_{i}})$ as well as corresponding text embeddings, $\mathbf{u_{i}}=f_{\phi}(\mathbf{w_{i}}).$ Then, the contrastive objective aims to match each frame with its corresponding utterance:
@@ 3.3 Variant 2: Language encoder via Transformer (CVCL+T)
In this variant, we examine the effect using a more powerful language encoder. In the previous variant, the language encoder treats the importance of each word in an utterance equally, regardless of its context. To test this, we replace the Embedding-based language encoder with a 2-layer Transformer Decoder architecture instead, enabling it to create contextualized word representations by dynamically weighting contributions from surrounding words. We refer to this variant as CVCL+T (where +T indicates the use of the Transformer-based language encoder).
Language Encoder (Transformer): We used a GPT-2 transformer decoder as our language encoder (Radford et al.,, 2019). Since GPT-2 uses a causal self-attention mask, the embedding from the last hidden state contains a contextualized aggregate representation of all of the preceding words in the utterance, allowing us to treat this representation as an utterance encoding that is further refined via contrastive learning with the vision encoder. All GPT-2 transformer models use 2 layers, 8 attention heads, and shared weights for the embedding and unembedding layers. We use the embedding of the last hidden state (corresponding to the <eos> token) as the output embedding, which is also of size $D=512$.
@@ 3.4 Variant 3: Language encoder via Transformer and language modeling
Language Encoder (Transformer): This variant utilizes the same 2-layer Transformer architecture as as Variant 2 as its language encoder. For the variant, we also separately utilize the decoder’s logits for predicting the next word in a given utterance.
Learning Objective (Contrastive with Language Modeling): In addition to the contrastive learning objective, for this variant, we incorporate an additional language modeling loss, similar to (Zhuang et al.,, 2024). The equation is given by:
@@
We also scale the contrastive loss relative to the language modeling loss by $\lambda_{C}$, and through a hyperparameter sweep, we set it to be 0.3 across all experiments (Zhuang et al.,, 2024).
@@ 3.5 Training Details
Each model was trained up to 100 epochs using the AdamW optimizer, with a learning rate of 1e-4 or 1e-5 depending on the model type. The learning rate schedule involved 5000 steps of linear warmup, followed by cosine annealing until the end of training. Early stopping was performed using the validation loss. [^3] To better track learning, we filter each validation set with CLIP to only contain frame-utterance pairs with high CLIP similarity scores. See Appendix B for additional details. For the vision encoder, we followed the specifications in Vong et al., (2024); Orhan and Lake, (2024): we used the the augmentations as in the vision encoder pretraining (described above), we sampled a randomly associated frame with each utterance, and we froze the vision encoder during training (except for the projection head). For the language encoder, we use a maximum sequence length of 48. To train the model, we use a fixed temperature parameter in the contrastive loss with $\tau=0.07$, with varying batch sizes depending on the model configuration. Vision dropout, language dropout and weight decay were all set to 0.1. We performed hyperparameter search over batch size, learning rate, vision dropout, language dropout, weight decay, and $\lambda_{c}$. For each dataset and model configuration, we trained 3 separate models with different random seeds. A full comparison between the various the architecture and training details is shown in Table 3.
@@ 4 Results
One of the main evaluation methods for the original CVCL model was examining its zero-shot classification performance, specifically using the Labeled-S evaluation dataset, as described in Section 2.2. This evaluation dataset consists of 22 different categories, with 100 4-way classification trials per category. Previously, Vong et al., (2024) showed that an Embedding-based multimodal neural network (CVCL) trained on the subset of manually transcribed data from baby S achieved 61.6% classification accuracy on this dataset. In this section, we compare and contrast the performance of the original CVCL model against additional variants of CVCL trained using Whisper-derived transcriptions.
@@
Figure 4: Labeled-S Classification Performance between CVCL trained with manual transcriptions (S-Manual-2022) vs. trained with Whisper transcriptions (S-Whisper-2022). Note that version of CVCL in the right plot utilizes a pre-trained vision transformer (ViT) instead of the CNN (ResNeXt), as well as a positional encoding scheme in the language encoder. Models were trained with three random seeds and results were averaged.
We first evaluated whether automatic transcriptions from Whisper would be comparable to manual transcriptions for training multimodal neural networks. These models were trained using a subset of Whisper transcribed utterances belonging to the same subset of videos that were previously manually transcribed and used to train CVCL, enabling a direct comparison of Whisper’s transcription quality. While initial analyses showed that the word frequency of the 22 categories from the Labeled-S evaluation showed a high correlation between the manually transcribed vs. Whisper transcriptions (Zhang et al.,, 2025), this did not provide any information about the quality of the Whisper-detected timestamps nor their temporal alignment with visual referents present in the dataset.
As can be seen in Figure 4, our results show that training CVCL on Whisper transcriptions provides comparable performance to CVCL, achieving a classification accuracy of 56.9%, compared to 61.6% from the original model. While there was a slight drop in performance compared to the original CVCL model, our results show that word-referent mappings are indeed learnable from Whisper-transcribed utterances, despite some of its limitations previously observed in transcription quality (Zhang et al.,, 2025). There are two additional reasons why the quality of Whisper transcriptions may have been sufficient to reach a similar level of performance as CVCL. First, rather than requiring perfect transcriptions for every word from every child-directed utterance, we only evaluate on a subset of the most prominent and commonly spoken nouns, for which there were a sufficient number of examples to learn from. Second, because of the relatively slow timescales at which visual information changes from the child’s perspective, large differences between the between relevant visual referents and detected utterance timings may not have been as important, something that was also observed with CVCL due to how some of its utterances were pre-processed during training.
@@ 4.2 Finding 2: Comparable performance for models trained on disjoint transcriptions
Figure 5: Labeled-S Classification Performance between CVCL trained on different combinations of Whisper transcriptions. The left-most bar (S-Whisper-2022, same as Figure 4), represents classification performance when trained on transcriptions the set of previously manually transcribed videos. The middle bar (S-Whisper-Disjoint) represents performance when trained only using transcriptions from new and non-overlapping videos from baby S. Finally, the right-most bar (S-Whisper) represents performance when trained using all of the combined Whisper-transcribed data from original and new videos. Models were trained with three random seeds and results were averaged.
@@
Next, to examine whether our model genuinely learns generalizable word-referent mappings, we conducted a follow-up experiment where the frames for training versus evaluation are completely disjoint. This contrasts with the previous experiment, where training and evaluation frames were independently extracted, but from the same set of videos. We trained a distinct set of models using the S-Whisper-Disjoint split, which utilizes Whisper transcriptions derived from an independent and non-overlapping set of videos from baby S, distinct from the videos used in both the S-Whisper-2022 videos and the Labeled-S evaluation set. Since all of the training data in this split comes from entirely new videos from baby S that were not previously manually transcribed, this enables us to examine whether the learned word-referent mappings generalizes to the Labeled-S dataset. In this instance, the frames from Labeled-S are now considered as new, previously unseen visual instances of these categories that were not seen at all during training for the S-Whisper-Disjoint models, and can be thought of as a form of cross-dataset generalization. This kind of exploration was not previously possible due to the limited available set of manually transcribed videos that were used for both training and evaluation (although the construction of each respective dataset was independent).
@@
The results are shown in Figure 5 (middle bar), and here we find that CVCL, when trained on non-overlapping data, also achieves a classification accuracy of 57.1%, indicating a relatively minimal drop in classification accuracy from generalizing within-distribution to generalizing out-of-distribution. Why do these models generalize well in this cross-dataset fashion? Despite being trained on a different set of videos, fundamentally, they are drawn from the same distribution of videos captured from a head-mounted camera in the same environment as the original manually transcribed videos for baby S, with similar objects and utterances across different videos. However, it is also the case that the contrastive objective forces multimodal representations to be robust and generalizable (Radford et al.,, 2019; Vong et al.,, 2024), guiding the models to learn the semantic relationship for different word-referent mappings, enabling broad generalization rather than overfitting or memorizing specific instances.
@@ 4.3 Finding 3: Classification performance increases with additional training data
Next, to determine whether increased training data could improve model performance, we trained models that used the maximum amount of training data available from baby S. These models were trained on the S-Whisper split, which was constructed using a combination of the Whisper transcriptions from S-Whisper-2022 and S-Whisper-Disjoint, resulting in approximately 3.5 times more data than used to train the original CVCL model (see Table 1 for additional details).
As shown in Figure 5 (right bar), we find that a CVCL model trained on this combined dataset split achieves an overall classification accuracy of 63.2%. While this outperforms CVCL, the gains in accuracy are relatively minimal given the large relative increase in training data. There are several possible explanations for the limited benefit to more training data. One reason similar to what was discussed in the previous section, the additional data from the newly transcribed videos in S-Whisper-Disjoint are being drawn from the same overall distribution of scenes and utterances, and therefore only providing a modest boost to the model’s generalization capabilities. An alternative explanation is that some categories from the Labeled-S evaluation set are sufficiently noisy or challenging, not just for CVCL or the variants we explore here, but for all models. For instance, Vong et al., (2024) evaluated CLIP ViT-L/14 on this dataset and found its zero-shot classification performance was only marginally higher (66.7%) despite being a strong vision-language foundation model, with CLIP also performing relatively poorly for a number of ambiguous categories like “floor” and “hand”.
@@ 4.4 Finding 4: Transformer-based language encoders achieve similar levels of classification performance
Next, to examine whether more sophisticated language encoders could improve grounded word learning, we replaced the Embedding-based language encoder with two variants involving a 2-layer Transformer. One of these variants used the same contrastive loss alone (CVCL+T), while the other variant combined this contrastive loss with an additional language modeling loss (CVCL+T+LM). These additional language encoder variants were chosen to examine whether other approaches for encoding utterances would be helpful for grounded word learning. The results in Figure 6 show that the Transformer-based language encoders show a similar or slightly worse performance compared to CVCL across the different splits, although no particular variant is consistently better or worse. One reason we may not have seen any additional advantages from stronger contextual-based encoding is the method of evaluation in our image classification tasks, which relies on only passing in the target category as a single token to the model during evaluation, rather than any additional words as context as seen during training.

### v2

@@ 3.1 Vision Encoder
For embedding static video frames, we replaced the ResNeXt-CNN in CVCL with a vision transformer (ViT-B/14) (Dosovitskiy et al., 2020), which is separately pre-trained only using visual data from the SAYCam dataset. This pre-training step is performed separately for each child’s visual data, using the self-supervised approach from DINO (Orhan and Lake, 2024; Caron et al., 2021). From the set of frames associated with a given utterance (up to 16 sampled frames), we always sample one frame randomly to be paired with the corresponding utterance. During training, data augmentation is performed by performing a random crop to each video frame, followed by resizing the frame to $224\times 224$ pixels, and occasionally horizontally flipping the frame. During evaluation, frames are instead always resized to 256 pixels along the minor edge, followed by a $224\times 224$ centered crop.
We obtain the output of the vision transformer from the [CLS] token, a single embedding of size 768. This embedding is fed through a learnable linear projection head, followed by a layer norm and a dropout layer, such that the vision encoder returns a single embedding of size $D=512$. Similar to Vong et al. (2024), the bulk of the vision encoder is frozen during training, except for the learnable linear projection head.
@@ 3.2 Variant 1: Language encoder with minimal syntax (CVCL)
In this model variant, we incorporate word-level positional embeddings into the language encoder to provide some sensitivity to word order, laying the groundwork for additional transformer-based variants described below. Because this is a relatively minor change from the original CVCL model, we still refer to this variant as CVCL throughout.
Language Encoder (Embedding): We used the same Embedding encoder from Vong et al. (2024), which extracts separate embeddings for each word in an utterance, and computes the average across all word embeddings to produce a single utterance embedding of size $D=512$. In this variant, we add a learned, absolute positional embedding to each token embedding before the averaging step. [^5] Because these learned positional embeddings are averaged across all words in the utterance, utterances of the same length will have the same positional embeddings, regardless of word order. We include positional embeddings for consistency with the transformer-based variants, which preserve finer-grained word-level positional information through the attention mechanism. Prior to passing utterances into the language encoder, tokenization was performed using a word-level tokenizer built via spaCy.
@@
Learning Objective (Contrastive): Following a standard approach to training vision-language models (Radford et al., 2021; Jia et al., 2021), we employ an InfoNCE contrastive learning objective, which aims to bring together frame embeddings with their corresponding utterance embeddings in a mini-batch (that is, encouraging a larger dot product between the vectors), while separating all other mismatching frame and utterance embeddings (smaller dot product between the vectors). First, we obtain visual embeddings for each frame, $\mathbf{v_{i}}=f_{\theta}(\mathbf{x_{i}})$ as well as corresponding text embeddings, $\mathbf{u_{i}}=f_{\phi}(\mathbf{w_{i}}).$ Then, the contrastive objective aims to match each frame with its corresponding utterance:
@@ 3.3 Variant 2: Language encoder via Transformer (CVCL+T)
In this variant, we examine the effect of using a more powerful language encoder. In the previous variant, the language encoder treats the importance of each word in an utterance equally, regardless of its context. To test this, we replace the Embedding-based language encoder with a 2-layer Transformer Decoder architecture instead, enabling it to create contextualized word representations by dynamically weighting contributions from surrounding words. We refer to this variant as CVCL+T (where +T indicates the use of the Transformer-based language encoder).
Language Encoder (Transformer): We used a GPT-2 transformer decoder as our language encoder (Radford et al., 2019). Since GPT-2 uses a causal self-attention mask, the embedding from the last hidden state contains a contextualized aggregate representation of all of the preceding words in the utterance, allowing us to treat this representation as an utterance encoding that is further refined via contrastive learning with the vision encoder. All GPT-2 transformer models use 2 layers, 8 attention heads. We use the embedding of the last hidden state (corresponding to the <eos> token) as the output embedding, which is also of size $D=512$.
@@ 3.4 Variant 3: Language encoder via Transformer and language modeling
Language Encoder (Transformer): This variant utilizes the same 2-layer Transformer architecture as Variant 2 for its language encoder. For the variant, we utilize the unembedding layer (which shares weights with the embedding layer) to generate the decoder’s logits for predicting the next word in a given utterance.
Learning Objective (Contrastive with Language Modeling): In addition to the contrastive learning objective, for this variant, we incorporate an additional language modeling loss, similar to (Zhuang et al., 2024a). The equation is given by:
@@
We scale the contrastive loss relative to the language modeling loss by $\lambda_{c}$, which controls the trade-off between vision-language alignment and language modeling. Following Zhuang et al. (2024a), we set $\lambda_{c}=0.3$ across all experiments with this variant, and validated this choice through a hyperparameter sweep. Since each child’s training data is shuffled and presented over multiple epochs, $\lambda_{c}$ represents a consistent weighting of these two objectives, independent of the child’s age in any given training example.
@@ 3.5 Training Details
Each model was trained up to 100 epochs using the AdamW optimizer, with a learning rate of 1e-4 or 1e-5 depending on the model type. The learning rate schedule involved 5000 steps of linear warmup, followed by cosine annealing until the end of training. Early stopping was performed using the validation loss. [^6] To better track learning, we filter each validation set with CLIP to only contain frame-utterance pairs with high CLIP similarity scores. See Appendix B for additional details. For the vision encoder, we followed the specifications in Vong et al. (2024); Orhan and Lake (2024): we used the augmentations as in the vision encoder pretraining (described above), we sampled a randomly associated frame with each utterance, and we froze the vision encoder during training (except for the projection head). For the language encoder, we use a maximum sequence length of 48. To train the model, we use a fixed temperature parameter in the contrastive loss with $\tau=0.07$, with varying batch sizes depending on the model configuration. Vision dropout, language dropout and weight decay were all set to 0.1. We performed hyperparameter search over batch size, learning rate, vision dropout, language dropout, weight decay, and $\lambda_{c}$. For each dataset and model configuration, we trained 3 separate models with different random seeds. A full comparison of the various architecture and training details is shown in Table 3.
@@ 4 Results
One of the main evaluation methods for the original CVCL model was examining its zero-shot classification performance, specifically using the Labeled-S evaluation dataset, as described in Section 2.2. This evaluation dataset consists of 22 different categories, with 100 4-way classification trials per category. Previously, Vong et al. (2024) showed that an Embedding-based multimodal neural network (CVCL) trained on the subset of manually transcribed data from baby S achieved 61.6% classification accuracy on this dataset. In this section, we compare and contrast the performance of the original CVCL model against additional variants of CVCL trained using Whisper-derived transcriptions.
@@
Figure 4: Labeled-S Classification Performance between CVCL trained with manual transcriptions (S-Manual-2022) vs. trained with Whisper transcriptions (S-Whisper-2022). Note that the version of CVCL in the right plot utilizes a pre-trained vision transformer (ViT) instead of the CNN (ResNeXt), as well as a positional encoding scheme in the language encoder. Models were trained with three random seeds. Error bars show bootstrapped 95% confidence intervals over category-level accuracies.
We first evaluated whether automatic transcriptions from Whisper would be comparable to manual transcriptions for training multimodal neural networks. These models were trained using a subset of Whisper transcribed utterances belonging to the same subset of videos that were previously manually transcribed and used to train CVCL, enabling a direct comparison of Whisper’s transcription quality. While initial analyses showed that the word frequency of the 22 categories from the Labeled-S evaluation showed a high correlation between the manually transcribed vs. Whisper transcriptions (Zhang et al., 2025), this did not provide any information about the quality of the Whisper-detected timestamps nor their temporal alignment with visual referents present in the dataset.
As can be seen in Figure 4, our results show that training CVCL on Whisper transcriptions provides comparable performance to CVCL, achieving a classification accuracy of 56.9%, compared to 61.6% from the original model. While there was a slight drop in performance compared to the original CVCL model, our results show that word-referent mappings are indeed learnable from Whisper-transcribed utterances, despite some of its limitations previously observed in transcription quality (Zhang et al., 2025). There are two additional reasons why the quality of Whisper transcriptions may have been sufficient to reach a similar level of performance as CVCL. First, rather than requiring perfect transcriptions for every word from every child-directed utterance, we only evaluate on a subset of the most prominent and commonly spoken nouns, for which there were a sufficient number of examples to learn from. Second, because of the relatively slow timescales at which visual information changes from the child’s perspective, large differences between relevant visual referents and detected utterance timings may not have been as important. This robustness to small mis-alignments in timing was observed with CVCL due to how some of its utterances were pre-processed during training.
@@ 4.2 Finding 2: Comparable performance for models trained on disjoint transcriptions
Figure 5: Labeled-S Classification Performance between CVCL trained on different combinations of Whisper transcriptions. The left-most bar (S-Whisper-2022, same as Figure 4), represents classification performance when trained on transcriptions from the set of previously manually transcribed videos. The middle bar (S-Whisper-Disjoint) represents performance when trained only using transcriptions from new and non-overlapping videos from baby S. Finally, the right-most bar (S-Whisper) represents performance when trained using all of the combined Whisper-transcribed data from original and new videos. Models were trained with three random seeds. Error bars show bootstrapped 95% confidence intervals over category-level accuracies.
@@
To examine whether our model genuinely learns generalizable word-referent mappings, we conducted a follow-up experiment where the frames for training versus evaluation are completely disjoint. This contrasts with the previous experiment, where training and evaluation frames were independently extracted, but from the same set of videos. We trained a distinct set of models using the S-Whisper-Disjoint split, which utilizes Whisper transcriptions derived from an independent and non-overlapping set of videos from baby S, distinct from the videos used in both the S-Whisper-2022 videos and the Labeled-S evaluation set. Since all of the training data in this split comes from entirely new videos from baby S that were not previously manually transcribed, this enables us to examine whether the learned word-referent mappings generalize to the Labeled-S dataset. In this instance, the frames from Labeled-S are now considered as new, previously unseen visual instances of these categories that were not seen at all during training for the S-Whisper-Disjoint models, and can be thought of as a form of within-child generalization. This kind of exploration was not previously possible due to the limited available set of manually transcribed videos that were used for both training and evaluation (although the construction of each respective dataset was independent).
@@
The results are shown in Figure 5 (middle bar), and here we find that CVCL, when trained on non-overlapping data, achieves a comparable classification accuracy of 57.1%, demonstrating robust within-child generalization to entirely new videos from the same child. Why do these models generalize well in this fashion? Despite being trained on a different set of videos, fundamentally, they are drawn from the same distribution of videos captured from a head-mounted camera in the same environment as the original manually transcribed videos for baby S, with similar objects and utterances across different videos. However, it is also the case that the contrastive objective forces multimodal representations to be robust and generalizable (Radford et al., 2019; Vong et al., 2024), guiding the models to learn the semantic relationship for different word-referent mappings, enabling broad generalization rather than overfitting or memorizing specific instances.
@@ 4.3 Finding 3: Classification performance increases with additional training data
To determine whether increased training data could improve model performance, we trained models that used the maximum amount of training data available from baby S. These models were trained on the S-Whisper split, which was constructed using a combination of the Whisper transcriptions from S-Whisper-2022 and S-Whisper-Disjoint, resulting in approximately 3.5 times more data than used to train the original CVCL model (see Table 1 for additional details).
As shown in Figure 5 (right bar), we find that a CVCL model trained on this combined dataset split achieves an overall classification accuracy of 63.2%. While this outperforms CVCL, the gains in accuracy are relatively minimal given the large relative increase in training data. There are several possible explanations for the limited benefit to more training data. One reason, similar to what was discussed in the previous section, is that the additional data from the newly transcribed videos in S-Whisper-Disjoint are being drawn from the same overall distribution of scenes and utterances, and therefore only providing a modest boost to the model’s generalization capabilities. An alternative explanation is that some categories from the Labeled-S evaluation set are sufficiently noisy or challenging, not just for CVCL or the variants we explore here, but for all models. For instance, Vong et al. (2024) evaluated CLIP ViT-L/14 on this dataset and found its zero-shot classification performance was only marginally higher (66.7%) despite being a strong vision-language foundation model, with CLIP also performing relatively poorly for a number of ambiguous categories like “floor” and “hand”.
@@ 4.4 Finding 4: Transformer-based language encoders achieve similar levels of classification performance
To examine whether more sophisticated language encoders could improve grounded word learning, we replaced the Embedding-based language encoder with two variants involving a 2-layer Transformer. One of these variants used the same contrastive loss alone (CVCL+T), while the other variant combined this contrastive loss with an additional language modeling loss (CVCL+T+LM). These additional language encoder variants were chosen to examine whether other approaches for encoding utterances would be helpful for grounded word learning. The results in Figure 6 show that the Transformer-based language encoders show a similar performance compared to CVCL across the different splits, although no particular variant is consistently better or worse. One reason we may not have seen any additional advantages is that our classification-based evaluation directly measures vision-language alignment, which all models are trained to optimize for via the contrastive objective. In a separate evaluation, we also examined word similarity across models (see Appendix C), finding that the models trained via a joint contrastive and language modeling objective demonstrated stronger correlations with human similarity judgments than a contrastive objective alone.